OpenStudy (anonymous):

FAN AND MEDAL HELPP?? What is the point-slope form of the equation of a line that passes through the point (3, 5) and has a slope of 2?

OpenStudy (anonymous):

y – y1 = m(x – x1) where m is the slope

OpenStudy (anonymous):

y - y1 = m(x - x1) where m is the slope

OpenStudy (anonymous):

thanks (: so witch one ?

OpenStudy (anonymous):

@khusker